What is the speed of 75 mm shaft transmitted 69.5 kW if stress is not to exceed 23.2 Mpa
Answer #1

Given:

  • Diameter of the shaft, D= 75 mm= 0.075 m
  • Power, P= 69.5 KW= 69.5×1000 W= 69500 W
  • Stress, 7 ​​​​​​​= 23.4 MPa= 23.4×10^6 Pa

To find:

  • Speed (N)

Note:

  • Polar moment of inertia of circular shaft, J= πD^4/32, where D= Diameter of the shaft in m
  • 1 Pa= 1 NN/m^2

Solution:

We know that, Power, P= 2πNT/60, where N= Speed in rpm, T= Torque in Nm.

We know that, by Torsion equation, Torque, T= 7 J/R, where 7 = Stress in N/m^2, J= Polar moment of inertia in m^4, R= Radius of the shaft in m

  • T=  \frac{\tau J}{R}
  • T=  \frac{\tau(\pi D^4/32) }{D/2}
  • T=. \frac{\tau(\pi D^3) }{16}
  • T= .\frac{23.2(10^6)(\pi)( 0.075^3) }{16}
  • T= 1921.771 Nm
  • Torque, T= 1921.771 Nm

​​​​​​​As said before, Power, P= 2πNT/60

  • P=  \frac{2\pi NT}{60}
  • 69500= \frac{2\pi N(1921.771)}{60}
  • N=  \frac{(69500)(60)}{2\pi (1921.771)}
  • N=  345.35 rpm
  • Speed, N= 345.35 rpm

​​​​​​​Answer:

  • Speed of the shaft= 345.35 rpm